AUL Test and electrical safety inspection:

Few minor issues were found. To be closed when recommendations are implemented.


Temporary solution  only accessible through the CERN GPN.
Original site ( ) stays deployed for the DIP/DIM live values, also accessible from outside through SSO.


GIF Logbook & elog  :   

Discussion ongoing. Not yet all needed features are implemented.
Not for this beam time.

Monitoring of primary lines to be done. This year we will have the monitoring of the neutral gas (Ar/CO2/Helium/N2), instead next year it will be done also for the other (iC4H10, R134a, SF6, CF4 etc.). This signals will be then available in PVSS.

On going discussion for main HFO gas supply line

EN-CV: New cooling line installed and operational in the preparation area.  

GIFCS-Pytimber app issue still under investigation


iC4H10 temporary line operational

-> issue due to the power cut and AUG test solved (plot attached)